If the y intercept of the graph 3x-ky=14 is -8 what Is the value of k

Answer by mananth(16949)

If the y intercept of the graph 3x-ky=14 is -8 what Is the value of k
3x-ky=14
..
ky = 3x-14
plug x=0
ky =-14
-14/k = y
-14/k = -8
k= -14/-8
k=7/4
..
m.ananth@hotmail.ca